# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 7.2.3. Graded multiplicities
Under the degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}_{\mathbb{C}}^{0}$, the modules $L(x\bullet_{l}0)_{\mathbb{C}}$, $M(x\bullet_{l}0)_{\mathbb{C}}$ and $Q(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}(x\bullet_{l}0)_{\mathbb{C}}$, $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}$ and $\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, such that the graded multiplicities are given by
(7.7) $(\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}:\widetilde{M}(y\bullet_{l}0)_{\mathbb{C}})_{\mathrm{gr}}=\begin{cases}q^{\prime}_{w_{0}y,w_{0}x}(v)&\text{if }y\bullet_{l}0\leq l\nu,\\ 0&\text{if else},\end{cases}$
@@
(7.8) $[\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}:\widetilde{L}(y\bullet_{l}0)_{\mathbb{C}}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),$
@@
Before proving Theorem 7.4, we firstly consider the graded multiplicities in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$. Set ${\mathrm{B}}$ as the $\Lambda$-graded (alternatively, $T$-equivariant) algebra with an isomorphism
$({\mathrm{B}})^{\mathrm{op}}=\bigoplus_{\lambda\in\Lambda}\mathrm{Hom}\big{(}\bigoplus_{w\in W}P(w\bullet 0)_{\mathbb{C}},\bigoplus_{w\in W}P(w\bullet 0-l\lambda)_{\mathbb{C}}\big{)}.$
@@
We denote by $L_{x}$, $Z_{x}$, $P_{x}$ the image of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$ ($x\in W_{\mathrm{ex}}$) under this equivalence. The following theorem is a recollection of results in [1, §18] (with conventions adopted from [39]).
@@
The liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ can be chosen such that ~$\widetilde{{\mathcal{L}}}_{t(\lambda)x}=\widetilde{{\mathcal{L}}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$), and
@@
Through the equivalences ${\mathrm{B}}\text{-}\mathrm{mod}^{T}\simeq u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$, the grading ${\mathrm{B}}^{\bullet}$ on ${\mathrm{B}}$ induces a different grading, hence a different graded multiplicity, on ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$. Our main task is to compare this graded multiplicity to the natural one induced by ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
By [12, §5.2.3, §5.5] (see also [11, Thm 4.5]), there is a $T\times{\mathbb{C}}^{\times}$-equivariant vector bundle ${\mathcal{E}}^{new}$ on $\tilde{{\mathfrak{g}}}$ satisfying Theorem 3.2(2), such that the graded algebra $\mathrm{End}^{{\mathrm{op}}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}}^{new})$ is Koszul (where grading is induced from ${\mathbb{C}}^{\times}$-equivariant structure). We can replace ${\mathcal{E}}$ by ${\mathcal{E}}^{new}$, then ${\mathrm{A}}^{\bullet}$ is a Koszul algebra. In particular, the graded algebra ${\mathrm{A}}_{0}^{\bullet}={\mathrm{A}}^{\bullet}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}$ satisfies: (1) ${\mathrm{A}}_{0}^{\bullet}$ is non-negatively graded; (2) ${\mathrm{A}}^{0}_{0}$ is a semi-simple algebra; (3) ${\mathrm{A}}_{0}^{\bullet}$ is generated by ${\mathrm{A}}^{1}_{0}$ over ${\mathrm{A}}^{0}_{0}$.
We can choose a lifting $\widetilde{{\mathcal{P}}}_{x}$ of ${\mathcal{P}}_{x}$ ($x\in W_{\mathrm{ex}}$) such that it is generated by degree 0 elements as a graded ${\mathrm{A}}_{0}^{\bullet}$-module. Define a graded algebra ${\mathrm{B}}^{\prime\bullet}$ via
@@
Then we have an isomorphism of algebras ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$. From the properties of ${\mathrm{A}}_{0}^{\bullet}$ (resp. Koszulity of ${\mathrm{B}}^{\bullet}$), we deduce that (1) ${\mathrm{B}}^{\prime\bullet}$ (resp. ${\mathrm{B}}^{\bullet}$) is non-negatively graded; (2) ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$) is a semi-simple algebra; (3) ${\mathrm{B}}^{\prime\bullet}$ (resp. ${\mathrm{B}}^{\bullet}$) is generated by ${\mathrm{B}}^{\prime 1}$ (resp. ${\mathrm{B}}^{1}$) over ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$). Hence by [6, Prop 2.4.1], we have $\mathrm{rad}({\mathrm{B}}^{\prime})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{\prime j}$ and $\mathrm{rad}({\mathrm{B}})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{j}$, and therefore we have a ($T$-equivariant) isomorphism of graded algebras
@@
It induces an equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}$. Note that the isomorphism (7.9) and the earlier isomorphism ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$ might be different, but they induce the same isomorphism on the semi-simple quotients ${\mathrm{B}}/\mathrm{rad}({\mathrm{B}})\simeq{\mathrm{B}}^{\prime}/\mathrm{rad}({\mathrm{B}}^{\prime})$. Therefore, under the composed equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, the simple module $\widetilde{L}_{x}$ is send to ~$\widetilde{{\mathcal{L}}}_{x}$ ($x\in W_{\mathrm{ex}}$) (up to shifts). As projective cover of $\widetilde{L}_{x}$, $\widetilde{P}_{x}$ is send to $\widetilde{{\mathcal{P}}}_{x}$ (up to shifts). Since $\widetilde{Z}_{x}$ is the projective cover of $\widetilde{L}_{x}$ in the Serre subcategory generated by $\widetilde{L}_{y}\langle d\rangle$ with $y\leq^{\frac{\infty}{2}}x$, it follows that $\widetilde{Z}_{x}$ is send to $\widetilde{{\mathcal{Z}}}_{x}$ (up to shifts).
@@
Proposition 7.6 might be proved as the analogue statement for restricted enveloping algebra in [11, Thm 1.2], instead of using Theorem 7.5.
@@
We fix the liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in Proposition 7.6. And choose liftings ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ such that ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{Z}}}_{x}$ and $\widetilde{{\mathcal{P}}}^{b}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{P}}}_{x}$. Then we have ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\lambda)x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{P}}}^{b}_{x}$). Apply $-\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ to any Verma flag of $\widetilde{{\mathcal{P}}}^{b}_{x}$, we obtain a Verma flag of $\widetilde{{\mathcal{P}}}_{x}$, since the corresponding statement in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ is true by the freeness of Verma modules over $Z_{\mathrm{Fr}}^{-}$. Hence we have
@@
Recall the $({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}=U{\mathfrak{n}}\otimes\widetilde{{\mathcal{P}}}^{b}_{x}$, which admits a Verma flag with factors of the form ${\mathbb{C}}_{\eta}\otimes\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)y}$, where $\eta\geq 0$ and ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}$ is a Verma factor of $\widetilde{{\mathcal{P}}}^{b}_{x}$. Moreover, the graded multiplicities satisfy the following equality in $\hat{{\mathbb{P}}}$,

### v4

@@ 7.2.3. Graded multiplicities
Let $\Omega$ be a bounded above poset ideal of $\Lambda$. Under the degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}_{\mathbb{C}}^{0}$, the modules $L(x\bullet_{l}0)_{\mathbb{C}}$, $M(x\bullet_{l}0)_{\mathbb{C}}$ and $Q(x\bullet_{l}0)_{\mathbb{C}}^{\Omega}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}(x\bullet_{l}0)_{\mathbb{C}}$, $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}$ and $\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\Omega}$ in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, such that the graded multiplicities are given by
[…]
@@
(7.10) $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}:\widetilde{L}(y\bullet_{l}0)_{\mathbb{C}}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),$
@@
Before proving Theorem 7.4, we firstly consider the graded multiplicities in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$. Set $\mathrm{B}$} as the $\Lambda$-graded (alternatively, $T$-equivariant) algebra with an isomorphism
${\mathrm{B}}^{\mathrm{op}}=\bigoplus_{\lambda\in\Lambda}\mathrm{Hom}\big(\bigoplus_{w\in W}P(w\bullet 0)_{\mathbb{C}},\bigoplus_{w\in W}P(w\bullet 0-l\lambda)_{\mathbb{C}}\big).$
@@
We denote by $L_{x}$, $Z_{x}$, $P_{x}$ the image of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$ ($x\in W_{\mathrm{ex}}$) under this equivalence. The following theorem is a recollection of results by Andersen–Jantzen–Soergel in [1, §18] (with conventions adopted from [37]).
@@
The liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ can be chosen such that $\widetilde{{\mathcal{L}}}_{t(\lambda)x}=\widetilde{{\mathcal{L}}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$), and
@@
Through the equivalences ${\mathrm{B}}\text{-}\mathrm{mod}^{T}\simeq u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}\simeq{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$, the grading $\mathrm{B}$$^{\bullet}$ on $\mathrm{B}$} induces a different grading, hence a different graded multiplicity, on ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$. Our main task is to compare this graded multiplicity to the natural one induced by ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
[…]
We can choose a lifting $\widetilde{{\mathcal{P}}}_{x}$ of ${\mathcal{P}}_{x}$ ($x\in W_{\mathrm{ex}}$) such that it is generated by degree 0 elements as a graded ${\mathsf{A}}_{0}^{\bullet}$-module. Define a graded algebra $\mathrm{B}$$^{\prime\bullet}$ via
@@
Then we have an isomorphism of algebras ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$. From the properties of ${\mathsf{A}}_{0}^{\bullet}$ (resp. Koszulity of $\mathrm{B}$$^{\bullet}$), we deduce that (1) $\mathrm{B}$$^{\prime\bullet}$ (resp. $\mathrm{B}$$^{\bullet}$) is non-negatively graded; (2) ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$) is a semi-simple algebra; (3) $\mathrm{B}$$^{\prime\bullet}$ (resp. $\mathrm{B}$$^{\bullet}$) is generated by ${\mathrm{B}}^{\prime 1}$ (resp. ${\mathrm{B}}^{1}$) over ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$). Hence by [5, Prop 2.4.1], we have $\mathrm{rad}({\mathrm{B}}^{\prime})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{\prime j}$ and $\mathrm{rad}({\mathrm{B}})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{j}$, and therefore we have a ($T$-equivariant) isomorphism of graded algebras
@@
It induces an equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}$. Note that the isomorphism (7.11) and the earlier isomorphism ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$ might be different, but they induce the same isomorphism on the semi-simple quotients ${\mathrm{B}}/\mathrm{rad}({\mathrm{B}})\simeq{\mathrm{B}}^{\prime}/\mathrm{rad}({\mathrm{B}}^{\prime})$. Therefore, under the composed equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, the simple module $\widetilde{L}_{x}$ is send to $\widetilde{{\mathcal{L}}}_{x}$ ($x\in W_{\mathrm{ex}}$) (up to shifts). As projective cover of $\widetilde{L}_{x}$, $\widetilde{P}_{x}$ is send to $\widetilde{{\mathcal{P}}}_{x}$ (up to shifts). Since $\widetilde{Z}_{x}$ is the projective cover of $\widetilde{L}_{x}$ in the Serre subcategory generated by $\widetilde{L}_{y}\langle d\rangle$ with $y\leq^{\frac{\infty}{2}}x$, it follows that $\widetilde{Z}_{x}$ is send to $\widetilde{{\mathcal{Z}}}_{x}$ (up to shifts).
@@
Proposition 7.6 might be proved as the analogue statement for restricted enveloping algebra in [9, Thm 1.2], instead of using Theorem 7.5.
@@
We fix the liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in Proposition 7.6. And choose liftings $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ such that $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{Z}}}_{x}$ and $\widetilde{{\mathcal{P}}}^{b}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{P}}}_{x}$. Then we have $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\lambda)x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{P}}}^{b}_{x}$).
Step 1. We firstly consider the multiplicity $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}$. Consider the following sequence induced by Koszul resolution
(7.12) $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}({\mathbb{C}}[{\mathfrak{n}}]\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow\widetilde{{\mathcal{Z}}}_{x}\rightarrow 0,$
is exact by the freeness of the $Z_{\mathrm{Fr}}^{-}$-module $M(x\bullet_{l}0)_{\mathbb{C}}$. By abuse of notation, we view $\langle\eta\rangle$ ($\eta\in\Lambda$) as a continuous ${\mathbb{Z}}[v^{\pm 1}]$-linear operator on $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ sending $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]$ to $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)x}]$, for any $x\in W_{\mathrm{ex}}$. By (7.12), we have the following equality in $K_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$,
(7.13) $\widetilde{{\mathcal{Z}}}_{x}]=\big(\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big)[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
In $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$, we express $\widetilde{{\mathcal{L}}}_{y}]$ into (topological) basis $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ and $\{[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]\}_{x}$ as
$\widetilde{{\mathcal{L}}}_{y}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{Z}}}_{x}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
Then (7.13) shows that
By [28, §11] (see also [36, Thm 6.1]), there is an equality
(7.15) $\sum_{x}(-1)^{\ell(x)+\ell(y)}q_{x,y}(v)\cdot p_{w_{0}x,w_{0}z}(v)=\delta_{y,z},\quad\forall y,z\in W_{\mathrm{ex}}.$
Hence it follows from Proposition 7.6 that
(7.16) $\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}=(-1)^{\ell(x)+\ell(y)}q_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Combining (7.1), (7.14) with (7.16), we deduce that
$\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}=(-1)^{\ell(x)+\ell(y)}p_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
By (7.15) again, we have
(7.17) $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Step 2. Now we study the multiplicity $(\widetilde{{\mathcal{Q}}}^{\Omega}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}$. Apply $\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ to any Verma flag of $\widetilde{{\mathcal{P}}}^{b}_{x}$, we obtain a Verma flag of $\widetilde{{\mathcal{P}}}_{x}$, since the corresponding statement in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ is true by the freeness of Verma modules over $Z_{\mathrm{Fr}}^{-}$. Hence we have
@@
Consider the $(T\times{\mathbb{C}}^{\times})$-equivariant ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module
$\widetilde{{\mathcal{Q}}}_{x}:=\mathrm{ind}_{c}(\widetilde{{\mathcal{P}}}^{b}_{x})=({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathsf{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x},$
[…]
(7.19) $0={\mathcal{Q}}_{0}\subset{\mathcal{Q}}_{1}\subset{\mathcal{Q}}_{2}\subset\cdots\subset{\mathcal{Q}}_{n}=\widetilde{{\mathcal{Q}}}_{x},$
with composition factors ${\mathcal{Q}}_{i+1}/{\mathcal{Q}}_{i}$ of the form $\mathrm{ind}_{c}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)$, where $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ are Verma factors of $\widetilde{{\mathcal{P}}}^{b}_{x}$. Since $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}$ is a ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module, there is an isomorphism of ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-modules $\mathrm{ind}_{c}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})\simeq\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}$, where $U{\mathfrak{n}}$ acts on $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}$ diagonally. There is a separated decreasing filtration
$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}={\mathcal{M}}_{0}\supset{\mathcal{M}}_{1}\supset\cdots,$
with composition factors ${\mathcal{M}}_{i}/{\mathcal{M}}_{i+1}$ of the form $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes{\mathbb{C}}_{\eta}$, $\eta\geq 0$. We formally write
$(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}:=\sum_{i=0}^{n-1}({\mathcal{Q}}_{i+1}/{\mathcal{Q}}_{i}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}.$
Then there is an equality in $\hat{{\mathbb{P}}}$,